Two cards are drawn without replacement from an ordinary deck of 52 cards. What are the odds in favor of drawing two queens?
A) 220:1
B) 1:220
C) 1:221
D) 1:168

Answer by edjones(8007)

nCr=n!/(n-r)!(r!) Combination of n things taken r at a time.
4C2/52C2 (2 of 4 queens)/(2 of 52 cards)
=6/1326
=1/221 Probability
=1-220 Odds in favor (B)
.
Ed